# Transportability of Prognostic Markers: Rethinking Common Practices through a Sufficient-Component-Cause Perspective

### v1

@@ Abstract
Transportability, the ability to maintain performance across populations, is a desirable property of of markers of clinical outcomes. However, empirical findings indicate that markers often exhibit varying performances across populations. For prognostic markers whose results are used to quantify of the risk of an outcome, oftentimes a form of updating is required when the marker is transported to populations with different disease prevalences. Here, we revisit transportability of prognostic markers through the lens of the foundational framework of sufficient component causes (SCC). We argue that transporting a marker “as is” implicitly assumes predictive values are transportable, whereas conventional prevalence-adjustment shifts the locus of transportability to accuracy metrics (sensitivity and specificity). Using a minimalist SCC framework that decomposes risk prediction into its causal constituents, we show that both approaches rely on strong assumptions about the stability of cause distributions across populations. A SCC framework instead invites making transparent assumptions about how different causes vary across populations, leading to different transportation methods. For example, in the absence of any external information other than disease prevalence, a cause-neutral perspective can assume all causes are responsible for change in prevalence, leading to a new form of marker transportation. Numerical experiments demonstrate that different transportability assumptions lead to varying degrees of information loss, depending on how population differ from each other in the distribution of causes. A SCC perspective challenges common assumptions and practices for marker transportability, and proposes transportation algorithms that reflect our knowledge or assumptions about how causes vary across populations.
@@
The underlying premise of reporting on the performance of biomarkers, tests, and prediction models (which we generally refer to as “markers”) is that performance metrics are assumed to be transportable from one population to another. In this context, we define transportability as conditions under which certain indices of marker performance remain the same in the new population as in the source population (where it was developed). Examples of such indices include accuracy (sensitivity and specificity), predictive values, discrimination, and calibration. However, in practice, this premise more often than not fails to hold. Research has repeatedly shown the performance of markers can vary significantly across populations, especially when the prevalence of the outcome of interest differs1. For example, risk prediction models for cardiovascular disease have shown substantial degradation in performance when transported to a new population, leading to substantial risk of harm (defined as the net benefit of using the model being lower than not using it)2. The likelihood of harm was reduced when models were updated to account for the difference in prevalences between their source population and the target population. A similar phenomenon was observed in risk prediction models for acute exacerbations of obstructive lung disease3. In another example, the performance of machine learning models for detecting pneumonia on chest X-rays substantially declined in data from settings not used to train the model4.
Often times, when transporting a marker for an outcome to a new population, some information about the characteristics of that population - colloquially referred to as ‘case-mix’- is available. One of the most common information about the new population is the outcome prevalence. For example, cancer registries often provide a good estimate of cancer risk in a population. This poses the challenge as to how such information can be used to revise our assessment of marker performance. For risk predictions models that return a quantitative estimate of disease (or outcome) risk given patient characteristics, some form of model updating (or revision) is often needed to correct for the under-estimation or over-estimation of risks in the new population5. When the model is a logistic regression, the most basic form of such updating involves modifying the intercept of the model to account for difference in prevalence between the source and target populations1,6,7. The underlying justification for this approach is that differences in outcome prevalence are mostly due to factors that act independently of predictors used in the model6. Because in a logistic regression model, changing the intercept is equivalent to applying a correcting odds-ratio to predicted risks, this method can be generalized to correcting the outputs of any algorithm that predicts the risk of a binary outcome, including black-box (e.g., machine learning) models8.
@@
Causal graphs are not the only model for causation. Another is the sufficient component causes (SCC) model14,15. SCC is a foundational model in that it establishes fully deterministic relationships between causes and effects (rather than a representation of statistical dependencies as in causal graphs). In this framework the variability in the outcome is due to the variability of, and interactions between, causes. This framework has recently been applied to explore the biologic plausibility of different link functions for modeling binary outcomes, resulting in proposals for more transportable measures of association such as the ‘switch relative risk’16 and more biology-aligned statistical models such as regression by composition17.
In this paper, we use a simple SCC model to study the most basic prediction setup: a binary factor that is used as a marker for the risk of an outcome. As a reference, we formulate this setup for prognostic markers, where adjustment for outcome prevalence seems to be a topical issue. We discuss how logical symmetries can be used to arrive at interpretations applicable to diagnostic markers. We create minimalist logic circuits to model a prognostic marker that provides partial, inconclusive information about an outcome of interest. This setup is the used to study marker transportability across populations, particularly as it relates to the variability in outcome prevalence. Our thesis is that by reducing the transportability problem to its basic constituents in this model, patterns will emerge that can provide insight into more complex scenarios.
@@
  • The latent binary variable $V$ represents all alternative causes - for example pathways related to the effect of tobacco smoking, which increases the risk of breast cancer even among those without a BRCA1 mutation. The absence of $V$ is responsible for false-negative marker values.
We assign the values of 1 and 0, respectively, to the ‘on’ and ‘off’ status of each of these switches. For the brevity of notations, by a simple character, we refer to the ‘on’ value, and by its dot-accented to its negated value (e.g., $\dot{U}=\neg U$ indicating $U=0$). We use the shorthand notation $P(X)$ for $P(X=1)$. For example, $P(T\dot{V})=P(T\dot{V}=1)=P(T=1,V=0)$.
@@
$D=(T\land U)\lor(V\land U)=(T\lor V)\land U.$
Before proceeding, we note that one can create an equally parsimonious SCC setup by swapping the logical AND and OR in the above equation, resulting in the setup $D=(T\land V)\lor U$. However, this is mathematically symmetrical to our reference setup. This becomes obvious by considering that the complementary marker (whose positive and negative results are swapped) is a marker for not experiencing the outcome under the reference setup: $\dot{D}=(\dot{T}\lor\dot{V})\land\dot{U}$. Because of this, any pattern we observe for PPV is also observed for NPV in this alternative formulation, albeit in the opposite direction, and vice versa. Similarly, any pattern we observe for SE in the reference setup is also observable, in the opposite direction, for SP in this alternative setup. Table 1 shows the minimum SCC setup, and the resulting 2-by-2 contingency tables of marker and outcome values ($P(T,D)$), from which performance metrics can be derived.
@@
Reference setup Reciprocal setup
@@
Given that $T$, $U$, and $V$ are all binary, a population is made up of 8 subgroups. For example, $T\dot{U}V$ is the subgroup where $T=1$, $U=0$, and $V=1$. As $U$ and $V$ are latent variables, the observed properties of the marker in a population is manifested in terms of $P(T,D)$, i.e., the two-by-two (contingency) table of marker by outcome status probabilities in that population.We express a contingency table by the sequence TP,FP,FN,TN{TP,FP,FN,TN}, where $TP=P(TD),FP=P(T\dot{D}),FN=P(\dot{T}D),TN=P(\dot{T}\dot{D})$.
If all we know about populations is their contingency tables, the full joint distribution $P(T,U,V)$ is unidentifiable, given its specification requires seven degrees of freedom but the contingency table only provides three. However, any joint distribution can be replaced by independent Bernoulli distributions for $P(T)$, $P(U)$, and $P(V)$ that would result in the same contingency table. Given specifying three Bernoulli distributions requires three degrees of freedom, this setup is identifiable and has a unique (1:1) mapping to a given contingency table. Because our focus is on metrics that are determined by the contingency table, without loss of generality the following we assume independence i.e., $P(T,U,V)=P(T)P(U)P(V)$.
Our interest is in the four fundamental performance characteristics of binary markers: sensitivity (SE=P(T|D)=TP/(TP+FN){SE}=P(T|D)=TP/(TP+FN)), specificity (${SP}=P(\dot{T}|\dot{D})=TN/(FP+TN)$), positive predictive value (PPV=P(D|T)=TP/(TP+FP){PPV}=P(D|T)=TP/(TP+FP)), and negative predictive value (${NPV}=P(\dot{T}|\dot{D})=TN/(FN+TN)$), as a function of prevalence ($P(D)=TP+FN$). By convention by SE and SP we refer to accuracy metrics, and by PPV and NPV as predictive values.
As a numerical example, consider a population where $P(T)=$ 0.25, $P(U)=$ 0.75, and $P(V)=$ 0.50. The 8 subgroups are scattered into the four cells of the contingency table (see Table 1). The contingency table emerging from this setup is {0.188,0.062,0.281,0.469}.This results in a prevalence ($P(D)$) of 0.469, SE of 0.400, SP of 0.882, PPV of 0.750, and NPV of 0.625. We will use this simple numerical example as a case study on various ways a marker can be transported.
@@
In this section we review the two conventional approaches for transporting a marker to a new population with potentially different outcome prevalence: transportation via accuracy (SE & SP), and transportation via predictive values (PPV & NPV).
In transportation by accuracy, we assume the SE and SP are equal between the source and target populations. The contingency table in the target population can be recovered by knowing outcome prevalence in the target population, denoted by $\pi$. For example, the proportion of true positives consists of those with the disease who also rest positive: $TP=\pi\times SE$, and so on. From this contingency table, various performance metrics can be derived. Details of this approach are provided in Appendix 1 (section A). This approach is the modus operandi for binary diagnostic markers, where the Bayes’ rule is used to combine pre-test probability with marker accuracy estimates to derive the post-test probability of the outcome21,22. Without any information that would distinguish the individual under evaluation, the pre-test probability is taken to be disease prevalence in the target population, which is equal to this mode of transportation.
@@
Transportation by predictive values refers to the where we assume it is the PPV and NPV from the source population that are transportable. Under this assumption, differences in outcome prevalence is due to the change in the prevalence of marker-positive individuals across populations. Again, if the target prevalence is known, PPV and NPV can be used to recover the contingency table in the target population. Details of this approach are provided in Appendix 1 (section B). This approach is more common for prognostic markers, a typical example being risk scoring tools that directly return an estimate of $P(D|T=x)$ for someone with marker value $x$. For a binary marker, this risk equation can be written as P(D|T=x)=(1NPV)+[PPV(1NPV)]xP(D|T=x)=(1-{NPV})+[{PPV}-(1-{NPV})]x. This equation indicates the implied assumption that the predictive values of the marker are transportable. In our numerical example, advertising prognostic information as a risk equation would result in $P(D|T)=0.375+0.375T$.
Returning to our numerical example, consider a population with 25% higher prevalence of breast cancer ($P(D)=0.586$). Transportation by accuracy results in a contingency table of {0.234,0.049,0.352,0.365}, corresponding to a PPV of 0.828 and NPV of 0.510. On the other hand, transportation by predictive values gives rise to a contingency table of {0.422,0.141,0.164,0.273}, corresponding to a SE of 0.720 and SP of 0.660.
Prevalance adjustment
A common way that prognostic markers are transported across populations is in the format of a risk equation for $P(D|x)$, either explicitly, as is the case in regression-based prediction models, or implicitly, as in black-box (e.g., machine learning) models. Consider applying the Bayes’ rule to the outcome risk conditional on marker value:
$P(D|T=x)=\frac{P(D)P(T=x|D)}{P(x)}=\frac{P(D)P(T=x|D)}{P(D)P(x|D)+P(\dot{D})P(x|\dot{D})}.$
@@
A familiar modeling framework for binary outcomes is the logistic regression. There, the logit function ($\mbox{logit}(x)=\log(\frac{x}{1-x})$) is used as link function connecting marker value to outcome probability. Applying this function to both sides of the risk equation for marker values, we have
$\mbox{logit}(P(D|T=x))=\mbox{logit}(P(D))+\log\left(\frac{P(T=x|D=1)}{P(T=x|D=0)}\right).$
@@
For binary markers, the LR is defined at two values. For $T=0$ it is (1SE)/SP(1-SE)/SP (aka negative LR), and for $T=1$ it is SE/(1SP)SE/(1-SP) (aka positive LR). For the marker to be transportable, both these LRs need to remain constant across populations, which will be the case if and only if both SE and SP remain constant. This in turn implies that prevalence adjustment for binary markers is equal to considering SE and SP to be transportable.
Under what conidtions are performance metrics transportable?
The above reasoning shows why adjustment for prevalence is generally expected to improve the transportability of a marker, as the conventional wisdom is that SE and SP, being defined within the diseased and non-diseased groups, are less dependent on prevalence23. While this might be intuitive as a general observation, under the SCC framework, none of these metrics are truly intrinsic. Rather, they emerge as properties of the distribution of causes, and their transportability depends on how this distribution varies across populations. One can indeed create population-generating mechanisms where a given subset of these metrics remains constant while others vary. However, the SCC framework helps us examine the plausibility of such mechanisms.
@@
  • As an example of derivations, consider PPV. Its transportability means $P(D|T)=c$ (a constant). But $P(D|T)=(P(TU\dot{V})+P(TUV))/P(T)=P(TU)/P(T)=P(U|T)$.
For the PPV to be transportable, the prevalence of universal causes among marker-positive individuals should remain constant. For NPV, transportability requires that among test-negative individuals, the proportion in whom both the universal and alternative causes are present should remain constant. These conditions are both satisfied if the prevalences of $U$ and $V$ are stable across populations. In this scenario, the entirety of variation in outcome prevalence is attributable to variation in $T$. In our breast cancer example, this would mean populations vary in breast cancer prevalence only because they differ in the BRCA1 mutation rate. For many markers, including BRCA1, this assumption is unrealistic, as many factors contribute to the variability in breast cancer risk24. In comparison, conditions for the transportation by accuracy are more complicated. It does not seem possible to specify a simple model on how $P(T,U,V)$ should vary such that both SE and SP remain stable. Such transportability method for a prognostic marker, as is implied in prevalence adjustment, amounts to a very specific set of assumptions on how causes vary across populations.
@@
  • The base setup is $P(T)=0.25$, $P(U)=0.75$, $P(V)=0.5$. Changes across populations are modeled by applying varying odds-ratios to the components that vary (labeled on top of each panel).
@@
The SSC framework enables us to express our beliefs on how causal pathways vary across populations, resulting in algorithms that are transparent and explicit in their underlying assumptions. Consider investigating a rare ancestral mutation as a prognostic marker for cancer that is known to have minimal variations across populations. In this case, it is reasonable to assume that $P(T)$ remains essentially constant across populations. When transporting this marker, our focus would therefore be on scenarios where differences in $P(U,V)$ derive variations in cancer prevalence. The resulting transportation method will be different from transportation by predictive values or accuracy. Indeed, both these transportation methods implicitly require that an increase in outcome prevalence be accompanied by change in marker positivity.
@@
But, what if the only information we have is differences in outcome prevalence across populations? We believe it is still worth visiting the plausibility of the assumptions under various transportability methods even under such a general case. For transportability by predictive values, one can question as to why only one cause ($T$, the one we happen to be measuring) should be responsible for change in prevalence. Transportability by SE and SP is also questionable due to unclear assumptions it places on the distribution of causes. For this general case, it might be more reasonable to take an explicit yet neutral stance about the variations in causes that bring about variations in prevalence, for example that all causes move by the same extent to cause change in prevalence - this approach makes no exception for thae causes for the mere fact that we are measuring it. As an implementation of this approach, one can solve for a common odds-ratio that, when applied to $P(T)$, $P(U)$, and $P(V)$ in the source population, results in a new population that matches the outcome prevalence in the target population. This “proportional odds” assumption is a new, distinct transportability approach that would generate different updated values of marker performance metrics compared with both conventional methods. Details of operationalizing this approach are provided in Appendix 1 (section C).
@@
In our numerical example, transporting the marker to a new population with 25% higher prevalence would require applying an odds-ratio of 1.434 to the three causes. This odd-ratio can be derived via standard univariate root finding (given that there is a monotonical relationship between this odds-ratio and prevalence). It can also be expressed as the real root of a cubic equation (see Appendix 1 - section C). The contingency table under this method of transportability will be {0.262,0.061,0.323,0.353}. This in turn results in SE of 0.448 and SP of 0.853. The PPV and NPV are, respectively, 0.811 and 0.522. Formulating marker performance as a risk equation would thus result in $P(D|T)=0.189+0.333T$). Note that both the intercept and slope (coefficient for $T$) of the risk equation are updated, which is different from the risk equation under the assumption of transportability of predictive values (where both remain unchanged), or accuracy (where only the intercept is updated).
@@
Ultimately, any assumption about how causes vary across populations will be a simplified version of a complex reality. The discrepancy between marker performance under a given transportability assumption and the true marker performance results in loss of prognostic information. We conducted brief simulation studies to explore such information loss under simple population-generating scenarios. The discrepancy between an assumed versus the true marker performance can be measured in different ways, each focusing on different consequences of such discrepancy (for example, through measures of discrimination, calibration, or prediction error). A more foundational approach is to measure information loss according to information theory, which in this context can be measured via the Kullback-Leibler divergence ($D_{KL}$)25. This information-theoretic measure quantifies the discrepancy between a true distribution and a candidate distribution in terms of ‘excess surprise’, which can be interpreted as the additional number of bits required to encode information from the true distribution using the candidate distribution, rather than using the true distribution itself26. In our case, these distributions are the true contingency table ($P(T,D)$) in the target population versus the one implied by a given transportability algorithm.

### v2

@@ Abstract
Transportability, the ability to maintain performance across populations, is a desirable property of of markers of clinical outcomes. However, empirical findings indicate that markers often exhibit varying performances across populations. For prognostic markers that are advertised as predictive risk equations, oftentimes a form of updating is required when the equation is transported to populations with different disease prevalences. Here, we revisit transportability of prognostic markers through the lens of the foundational framework of sufficient component causes (SCC). We argue that transporting a marker “as is” implicitly assumes predictive values are transportable, whereas conventional prevalence adjustment shifts the locus of transportability to accuracy metrics (sensitivity and specificity). Using a minimalist SCC framework that decomposes risk prediction into its causal constituents, we show that both approaches rely on strong assumptions about the stability of cause distributions. A SCC framework instead invites making transparent assumptions about how different causes vary across populations, leading to different transportation methods. For example, in the absence of any external information other than disease prevalence, a cause-neutral perspective can assume all causes are responsible for change in prevalence, leading to a new form of marker transportation. Numerical experiments demonstrate that different transportability assumptions lead to varying degrees of information loss, depending on the distribution of causes. A SCC perspective challenges common assumptions and practices for marker transportability, and proposes transportation algorithms that reflect our knowledge or assumptions about how causes vary across populations.
@@
The underlying premise of reporting on the performance of biomarkers, tests, and prediction models (which we generally refer to as “markers”) is that performance metrics are transportable from one population to another. In this context, we define transportability as conditions under which certain indices of marker performance remain the same across populations. Examples of such indices include accuracy (sensitivity and specificity), predictive values, discrimination, and calibration. However, in practice, this premise more often than not fails to hold. Research has repeatedly shown the performance of markers can vary significantly across populations, especially when the prevalence of the outcome of interest differs1. For example, risk prediction models for cardiovascular disease have shown substantial degradation in performance when transported to a new population, leading to substantial risk of harm (defined as the net benefit of using the model being lower than not using it)2. The likelihood of harm was reduced when models were updated to account for the difference in prevalences between their source population and the target population. A similar phenomenon was observed in risk prediction models for acute exacerbations of obstructive lung disease3. In another example, the performance of machine learning models for detecting pneumonia on chest X-rays substantially declined in data from settings not used to train the model4.
Often times, when transporting a marker for an outcome to a new population, some information about the characteristics of that population - colloquially referred to as case-mix- is available. One of the most common information pieces is the outcome prevalence. For example, cancer registries often provide a good estimate of cancer risk in a population. This poses the challenge as to how such information can be used to revise our assessment of marker performance when transporting it to a new population. For risk predictions models that return a quantitative estimate of disease (or outcome) risk given patient characteristics, some form of model updating (or revision) is often needed to correct for the under-estimation or over-estimation of risks in the new population5. When the model is a logistic regression, the most basic form of such updating involves modifying the intercept of the model to account for difference in prevalence between the source and target populations1,6,7. The underlying justification for this approach is that differences in outcome prevalence are mostly due to factors that are distributed, and act, independently of predictors used in the model6. Because in a logistic regression model, changing the intercept is equivalent to applying an odds-ratio to predicted risks, this method can be generalized to applying a correcting odds-ratio to the outputs of any algorithm that predicts the risk of a binary outcome, including black-box (e.g., machine learning) models8.
@@
Causal graphs are not the only model for causation. Another is the sufficient component causes (SCC) model14,15. SCC is a foundational model in that it establishes fully deterministic relationships between causes and effects (rather than a representation of statistical dependencies as in causal graphs). This framework has recently been applied to explore the biologic plausibility of different link functions for modeling binary outcomes, resulting in proposals for more transportable measures of association such as the ‘switch relative risk’16 and more biology-aligned statistical models such as regression by composition17.
In this paper, we use a simple SCC model to study the most basic prediction setup: a binary factor that is used as a marker for the risk of a binary outcome. As a reference, we formulate this setup for prognostic markers, where adjustment for outcome prevalence seems to be a topical issue. We later discuss how symmetries can be used to investigate the transportability of diagnostic markers but leave deeper investigations to future work. This setup is the used to study marker transportability across populations, particularly as it relates to the variability in outcome prevalence. Our thesis is that by reducing the transportability problem to its basic constituents in this model, patterns will emerge that can provide insight into more complex scenarios.
@@
  • The latent binary variable $V$ represents all alternative causes - for example pathways related to the effect of tobacco smoking, which increases the risk of breast cancer even among those without a BRCA1 mutation. The presence of $V$ is responsible for false-negative marker values.
We assign the values of 1 and 0, respectively, to the ‘on’ and ‘off’ status of each of these switches. For the brevity of notations, by a simple character, we refer to the ‘on’ value, and by its dot-accented to its negated value (e.g., $\dot{U}=\neg U$ indicating $U=0$).
@@
$D=(T\land U)\lor(V\land U)=(T\lor V)\land U,$
where $\land$ and $\lor$ are logical AND and OR, respectively.
Given that $T$, $U$, and $V$ are all binary, a population is made up of 8 subgroups. As $U$ and $V$ are latent variables, the observed properties of the marker in a population is manifested in terms of $P(T,D)$, i.e., the two-by-two (contingency) table of marker by outcome status probabilities. We represent the contingency table by the sequence TP,FP,FN,TN{TP,FP,FN,TN}, where $TP=P(TD=1),FP=P(T\dot{D}=1),FN=P(\dot{T}D=1),TN=P(\dot{T}\dot{D}=1)$ are, respectively, true positive, false positive, false negative, and true negative probabilities.
Our interest is in the four fundamental performance characteristics of binary markers: positive predictive value (PPV=P(D=1|T=1)=TP/(TP+FP){PPV}=P(D=1|T=1)=TP/(TP+FP)), negative predictive value (NPV=P(T=1|D=0)=TN/(FN+TN){NPV}=P(T=1|D=0)=TN/(FN+TN)), sensitivity (SE=P(T=1|D=1)=TP/(TP+FN){SE}=P(T=1|D=1)=TP/(TP+FN)), and specificity (SP=P(T=0|D=0)=TN/(FP+TN){SP}=P(T=0|D=0)=TN/(FP+TN)), as a function of prevalence ($P(D=1)=TP+FN$). By convention PPV and NPV are referred to as predictive values and SE and SP as accuracy metrics.
Before proceeding, we note that one can create an equally parsimonious SCC setup by swapping the logical AND and OR in the above equation, resulting in the setup $D=(T\land V)\lor U$. However, this is mathematically symmetrical to our reference setup. This becomes obvious by considering that the complementary marker (whose positive and negative results are swapped) is a marker for not experiencing the outcome under the reference setup: $\dot{D}=(\dot{T}\lor\dot{V})\land\dot{U}$. Because of this, any pattern we observe for PPV in the main formulation is also observed for NPV in this alternative formulation, albeit in the opposite direction, and vice versa. Similarly, any pattern we observe for SE in the reference setup is also observable, in the opposite direction, for SP in this alternative setup. Table 1 shows the minimum SCC setup, and the resulting 2-by-2 contingency tables of marker and outcome values ($P(T,D)$), from which performance metrics can be derived.
@@
Reference setup Symmetrical setup
@@
  • Subgroups are defined as the combination of causes such that the product evaluates to 1. For example, $T\dot{U}V$ is the subgroup where $T=1$, $U=0$, and $V=1$.
Mapping between contingency table and cause probabilities
If all we know about marker performance in a populations is its contingency table, the full joint distribution $P(T,U,V)$ is unidentifiable, given its specification requires seven degrees of freedom but the contingency table only provides three. However, any joint distribution can be replaced by independent marginal distributions for $P(T=1)$, $P(U=1)$, and $P(V=1)$ that would result in the same contingency table. We use the shorthand notations $P_{T}=P(T=1),P_{U}=P(U=1)$, and $P_{V}=P(V=1)$. Given specifying three Bernoulli distributions for these probabilities requires three degrees of freedom, this setup is identifiable and has a unique (1:1) mapping to a given contingency table. This 1:1 mapping can thus be formulated as follows:
  • From {TP,FP,FN,TNTP,FP,FN,TN} to $P_{T},P_{U},P_{V}$}:
    $P_{T}=TP+FP,\quad P_{U}=TP/(TP+FP),\quad P_{V}=FN/[(1-P_{T})P_{U}]$
  • From $P_{T},P_{U},P_{V}$} to {TP,FP,FN,TNTP,FP,FN,TN}:
    $TP=P_{T}P_{U},\quad FP=P_{T}(1-P_{U}),\quad FN=(1-P_{T})P_{U}P_{V},\quad TN=1-TP-FP-FN.$
As a numerical example, consider a population where $P_{T}=$ 0.25, $P_{U}=$ 0.75, and $P_{V}=$ 0.50. The 8 subgroups are scattered into the four cells of the contingency table (see Table 1), which is presented in the top row of Table 2. This results in a prevalence of 0.469, SE of 0.400, SP of 0.882, PPV of 0.750, and NPV of 0.625. We will use this simple numerical example as a case study on various ways a marker can be transported.
Table 2: Numerical example of different marker transportation methodology
Description { $P_{T},P_{U},P_{V}$ } Contingency table SE SP PPV NPV
Source population {0.250,0.750,0.500} {0.188,0.062,0.281,0.469} 0.400 0.882 0.750 0.625
Target population {0.333,0.800,0.667} {0.267,0.067,0.356,0.311} 0.429 0.824 0.800 0.467
By predictive values {0.333,0.750,0.500} {0.250,0.083,0.250,0.417} 0.500 0.833 0.750 0.625
By accuracy {0.293,0.848,0.623} {0.249,0.044,0.373,0.333} 0.400 0.882 0.848 0.472
Proportional odds {0.349,0.829,0.617} {0.290,0.060,0.333,0.318} 0.465 0.841 0.829 0.489
  • Contingency table presents the sequence TP (true positive), FP (false positive), FN (false negative), TN (true negative)
@@
In this section we review the two conventional approaches for transporting a marker to a new population with potentially different outcome prevalence: transportation via accuracy (SE & SP), and transportation via predictive values (PPV & NPV). Ultimately, using a marker for decision-making entails interpreting its predictive values - i.e., $P(D|T)$. Different transportation methods construct these predictive values by combining different pieces of information from the source and target populations.
Transportation by predicted values
@@
Transportation by predictive values refers to the where predictive values from the source population are considered to be the same as in the target population. Given that $P(T,D)=P(T)P(D|T)$, this transportability means our estimate of the contingency table in the target population is determined by $P(T)$ in that population, combined with predictive values from the source population. This approach is more common for prognostic markers, a typical example being risk scoring tools that directly return an estimate of $P(D=1|T)$ for someone with marker value $x$. For a binary marker, this risk equation can be written as P(D=1|T)=(1NPV)+[PPV(1NPV)]TP(D=1|T)=(1-{NPV})+[{PPV}-(1-{NPV})]T. In our numerical example, advertising prognostic information as a risk equation would result in $P(D=1|T)=0.375+0.375T$.
Transportation by accuracy
In transportation by accuracy, we assume that $P(T|D)$, i.e., the SE and SP, are equal between the source and target populations. In order to construct predictive probabilities, we apply the Bayes’ rule:
$P(D=1|T)=\frac{P(D=1)P(T|D=1)}{P(T)}=\frac{P(D=1)P(T|D=1)}{P(D=1)P(T|D=1)+P(D=0)P(T|D=0)}.$
Thus, under this transportability scheme we need access to prevalence in the target population to derive predictive values. This approach is the modus operandi for binary diagnostic markers, where the Bayes’ rule is used to combine pre-test probability with marker accuracy estimates to derive the post-test probability of the outcome21,22. Without any information that would distinguish the individual under evaluation, the pre-test probability is taken to be disease prevalence in the target population, which is equal to this mode of transportation.
Returning to our running example, imagine we are transporting the above-mentioned prognostic marker to a population where $P_{T}=$ 0.250, $P_{U}=$ 0.750, $P_{V}=$ 0.500. The true contingency table and marker performance metrics are provided in Table 2. Transportation by predictive values preserves PPV and NPV in the target population, but will result in SE and SP of 0.5 and 0.833. Now, imagine we know the target population prevalence, which is 0.622 (32.7% higher than in the source population). If we consider transportation by accuracy, we will arrive at the PPV and NPV values of, respectively, 0.750 and 0.625.
Prevalence adjustment for risk equations
@@
A common way that prognostic markers are transported across populations is in the format of a risk equation for $P(D|T)$, either explicitly, as is the case in regression-based prediction models, or implicitly, as in black-box (e.g., machine learning) models. A familiar modeling framework for binary outcomes is the logistic regression. There, the logit function ($\mbox{logit}(x)=\log(\frac{x}{1-x})$) is used as link function connecting marker value to outcome probability. Applying this function to both sides of the previous equation, we have
$\mbox{logit}(P(D=1|T))=\mbox{logit}(P(D=1))+\log\left(\frac{P(T|D=1)}{P(T|D=0)}\right).$
@@
For binary markers, the LR is defined at two values. For $T=0$ it is (1SE)/SP(1-SE)/SP (aka negative LR), and for $T=1$ it is SE/(1SP)SE/(1-SP) (aka positive LR). For the marker to be transportable, both these LRs need to remain constant across populations, which will be the case if and only if both SE and SP remain constant. This in turn implies that prevalence adjustment for binary markers is equal to transportation by accuracy; i.e., considering SE and SP to be transportable.
Under what conditions are performance metrics transportable?
The above reasoning shows why adjustment for prevalence is generally expected to improve transportability, as the conventional wisdom is that SE and SP, being defined within the diseased and non-diseased groups, are less dependent on prevalence23. While this might be intuitive as a general observation, under the SCC framework, none of these metrics are truly intrinsic. Rather, they emerge as properties of the distribution of causes, and their transportability depends on how this distribution varies across populations. One can indeed create population-generating mechanisms where a given subset of these metrics remains constant while others vary. However, the SCC framework helps us examine the plausibility of such mechanisms.
@@
  • As an example of derivations, consider PPV. Its transportability means $P(D=1|T=1)=c$ (a constant). But $P(D=1|T=1)=(P(TU\dot{V}=1)+P(TUV=1))/P(T=1)=P(TU=1)/P(T=1)=P(U=1|T=1)$.
For the PPV to be transportable, the prevalence of universal causes among marker-positive individuals should remain constant. For NPV, transportability requires that among test-negative individuals, the proportion in whom both the universal and alternative causes are present should remain constant. These conditions are both satisfied if the prevalences of $U$ and $V$ are stable across populations. In this scenario, the entirety of variation in outcome prevalence is attributable to variation in $T$. In our breast cancer example, this would mean populations vary in breast cancer prevalence only because they differ in the BRCA1 mutation rate. For many outcomes, including breast cancer, this assumption is unrealistic, as many factors contribute to the variability in breast cancer risk24. In comparison, conditions for the transportation by accuracy are more complicated. It does not seem possible to specify a simple model on how $P(T,U,V)$ should vary such that both SE and SP remain stable. Such transportability method for a prognostic marker, as is implied in prevalence adjustment, amounts to a very specific set of assumptions on how causes vary across populations.
@@
  • The base setup is $P_{T}=0.25$, $P_{U}=0.75$, $P_{V}=0.5$. Changes across populations are modeled by applying varying odds-ratios to the components that vary (labeled on top of each panel).
@@
The SSC framework enables us to express our beliefs on how causal pathways vary across populations, resulting in algorithms that are transparent and explicit in their underlying assumptions. Consider investigating a rare ancestral mutation as a prognostic marker for cancer that is known to have minimal variations across populations. In this case, it is reasonable to assume that $P_{T}$ remains essentially constant across populations. When transporting this marker, our focus would therefore be on scenarios where differences in $P(U,V)$ derive variations in cancer prevalence. The resulting transportation method will be different from transportation by predictive values or accuracy. Indeed, both these transportation methods implicitly require that an increase in outcome prevalence be accompanied by change in marker positivity.
Proportional-odds transportability method
@@
But, what if the only information we have is differences in outcome prevalence across populations? We believe it is still worth visiting the plausibility of the assumptions under various transportability methods even under such a general case. For transportability by predictive values, one can question as to why only one cause ($T$, the one we happen to be measuring) should be responsible for change in prevalence. Transportability by accuracy is also questionable due to unclear assumptions it places on the distribution of causes. For this general case, it might be more reasonable to take a neutral yet transparent stance about the variations in causes, for example that all causes move by the same extent to cause change in prevalence. As an implementation of this approach, instead of applying a correcting odds-ratio to predicted values, one can solve for a common odds-ratio that, when applied to $P_{T}$, $P_{U}$, and $P_{V}$ in the source population, results in a new population that matches the outcome prevalence in the target population. This “proportional odds” assumption is a new, distinct transportability approach that would generate different updated values of marker performance metrics compared with both conventional methods. Details of operationalizing this approach are provided in Appendix 2.
@@
In our numerical example (last row of Table 2), we need to apply an odds-ratio of 1.612 to the three causes to match the target population’s outcome prevalence. This odd-ratio can be derived via standard univariate root finding - given that there is a monotonical relationship between this odds-ratio and prevalence. It can also be expressed as the real root of a cubic equation for which standard solutions are available (see Appendix 2). The contingency table under this method of transportability will be {0.290, 0.060, 0.333, 0.318}. This in turn results in SE of 0.465 and SP of 0.841. The PPV and NPV are, respectively, 0.829 and 0.489. Formulating marker performance as a risk equation would thus result in $P(D|T)=0.171+0.317x$. Note that both the intercept and slope (coefficient for $T$) of the risk equation are updated, which is different from the risk equation under the assumption of transportability of predictive values (where both remain unchanged), or accuracy (where only the intercept is updated).
@@
Ultimately, any assumption about how causes vary across populations will be a simplified version of a complex reality. The discrepancy between marker performance under a given transportability assumption and the true marker performance results in loss of prognostic information. We conducted brief simulation studies to explore such information loss under simple population-generating scenarios. The discrepancy between an assumed versus the true marker performance can be measured in different ways, each focusing on different consequences of such discrepancy (for example, through measures of discrimination, calibration, or prediction error). A more foundational approach is to measure information loss according to information theory. In this context, the relevant measure would be the Kullback-Leibler divergence ($D_{KL}$)25. This information-theoretic measure quantifies the discrepancy between a true distribution and a candidate distribution in terms of ‘excess surprise’, which can be interpreted as the additional number of bits required to encode information from the true distribution using the candidate distribution, rather than using the true distribution itself26. In our case, these distributions are the true contingency table ($P(T,D)$) in the target population versus the one implied by a given transportability algorithm.